Help with 1-5 please

Help with 1-5 please
Answer #1
Journal entry
Date General Journal Debit Credit
1/1/2019 Cash 3,456,448
discount on bonds 543,552
bonds payable 4,000,000
2-a) par maturity value Annual rate / year semi annual cash payment
4,000,000 * 6% 6./12 120000
semi annual Straight line
2-b) par value bonds price Discount periods disc amortization
4,000,000 - 3,456,448 = 543,552 / 30 = 18118
2-c) Semi annual cash Discount bond interest expense
payment amortization
120,000 + 18118 = 138,118
3) total bond interest expense over life of bonds
amount repaid
30 payments of 120,000 3600000
par value ant maturity 4,000,000
total repaid 7600000
less amount borrowed 3,456,448
total bond interest expense. 4,143,552
4) unamort Carrying
period discount value
1/1/2019 543,552 3,456,448
6/30/2019 525,434 3,474,566
12/31/2019 507,315 3,492,685
6/30/2020 489,197 3,510,803
12/31/2020 471,078 3,528,920
5)
Date General Journal Debit Credit
6/30/2019 interest expense 138,118
Discount on bonds payable 18,118
cash 120,000
31/12/2019
interest expense 138,118
discount on bonds payable 18,118
cash 120,000